# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products

### v2

@@ 3.3.2 Proposed residual-based Chebyshev filtering approach for generalized eigenproblem
$\leq\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\tilde{\eta}_{r}\tilde{\zeta}\lVert\textbf{R}^{(i)}\rVert\,+\left\lvert b_{r}\right\rvert\,\tilde{\eta}_{r}\tilde{\zeta}\lVert\textbf{R}^{(i)}\rVert\,+\lvert c_{r}\rvert\,\tilde{\eta}_{r-1}\tilde{\zeta}\lVert\textbf{R}^{(i)}\rVert\,+\left\lvert a_{r}\right\rvert\,\tilde{\zeta}\lVert\textbf{R}^{(i)}\bm{\Lambda}_{r}^{(i)}+\textbf{A}\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{r}^{(i)}\rVert\,$
$\leq\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\tilde{\eta}_{r}\tilde{\zeta}\lVert\textbf{R}^{(i)}\rVert\,+\left\lvert b_{r}\right\rvert\,\tilde{\eta}_{r}\tilde{\zeta}\lVert\textbf{R}^{(i)}\rVert\,+\lvert c_{r}\rvert\,\tilde{\eta}_{r-1}\tilde{\zeta}\lVert\textbf{R}^{(i)}\rVert\,+\left\lvert a_{r}\right\rvert\,\tilde{\zeta}\left(\lVert\bm{\Lambda}_{r}^{(i)}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,+\lVert\textbf{A}\textbf{D}^{-1}\textbf{B}\rVert\,\lVert\underline{\textbf{R}}^{(i)}_{r}\rVert\,\right)$
We further note that using Lemma 3.2 we can write $\lVert\underline{\textbf{R}}_{r}^{(i)}\rVert\,=\lVert\bm{\Delta}_{r}^{(i)}+\textbf{R}_{r}^{(i)}\rVert\,\leq\tilde{\zeta}\tilde{\eta}_{r}\lVert\textbf{R}^{(i)}\rVert\,+\lVert\textbf{R}_{r}^{(i)}\rVert\,\leq\tilde{\zeta}\tilde{\eta}_{r}\lVert\textbf{R}^{(i)}\rVert\,+f_{r}\lVert\textbf{R}^{(i)}\rVert\,$, and consequently we can write
$\lVert\bm{\Delta}_{r+1}^{(i)}\rVert\,\leq\Biggl{[}\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\tilde{\eta}_{r}+\left\lvert b_{r}\right\rvert\,\tilde{\eta}_{r}+\lvert c_{r}\rvert\,\tilde{\eta}_{k-1}+\left\lvert a_{r}\right\rvert\,(\lVert\bm{\Lambda}_{r}^{(i)}\rVert\,+\lVert\textbf{A}\textbf{D}^{-1}\textbf{B}\rVert\,(\tilde{\zeta}\tilde{\eta}_{r}+f_{k}))\Biggr{]}\tilde{\zeta}\lVert\textbf{R}^{(i)}\rVert\,$ (47)
Defining $\tilde{\eta}_{r+1}=\left\lvert a_{r}\right\rvert\,\lVert\textbf{H}\rVert\,\tilde{\eta}_{r}+\left\lvert b_{r}\right\rvert\,\tilde{\eta}_{r}+\lvert c_{r}\rvert\,\tilde{\eta}_{r-1}+\left\lvert a_{r}\right\rvert\,(\lVert\bm{\Lambda}_{r}^{(i)}\rVert\,+\lVert\textbf{A}\textbf{D}^{-1}\textbf{B}\rVert\,(\tilde{\zeta}\tilde{\eta}_{r}+f_{r}))$. Hence, using induction, we argue that $\lVert\bm{\Delta}_{k+1}^{(i)}\rVert\,\leq\tilde{\zeta}\tilde{\eta}_{k+1}\lVert\textbf{R}^{(i)}\rVert\,$ for $k=0,1,\dots,p$. ∎
@@
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\tilde{\zeta}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (48)
If this inequality is satisfied for at $i=i_{0}$ then it holds for all $i>i_{0}$
Proof.
As a consequence of Lemmas 3.2, 3.5 and 3.4 we have $\lVert\hat{\bm{\Delta}}_{p}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,\tilde{\zeta}\tilde{\eta}_{p}\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ and in order to have $\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})<\angle(\mathcal{S}^{(i)},\mathcal{S})$ from Eq. 4 we require
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\tilde{\zeta}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (49)
@@ 4 Results and Discussion
In the following subsections, we focus on specific problem classes, starting with the results for standard eigenvalue problems, including the cases of real symmetric and complex Hermitian matrices, followed by generalized eigenvalue problems, where we assess the convergence behavior and the performance of the ChFSI and the R-ChFSI methods. To this end, we consider a set of three benchmark systems whose dimensions are summarized in Table 1. These eigenvalue problems are obtained from the finite-element discretization of the Kohn-Sham DFT equations for the periodic supercells of various sizes for body-centered cubic (BCC) Molybdenum systems with a single vacancy. For all the benchmarks we utilize a Chebyshev polynomial degree of 75.
@@
Table 1: Dimensions of the benchmark problems considered. These correspond to 3 material systems comprising $6\times 6\times 6$, $8\times 8\times 8$ and $10\times 10\times 10$ body-centered-cubic (BCC) periodic supercells of Molybdenum with a single vacancy
@@ 4.1 Standard Eigenvalue Problems
We now evaluate the accuracy and performance of the proposed R-ChFSI method as described in Algorithm 2 for solving the standard eigenvalue problems. We choose the eigenproblem encountered in using the higher-order finite-element discretization of Kohn-Sham DFT as outlined in Das et al. [38], Motamarri et al. [55]. We begin by comparing the R-ChFSI method to the ChFSI method using FP32 and TF32 arithmetic. We demonstrate below that the R-ChFSI method with FP32 and TF32 arithmetic converges to the desired residual tolerance in a similar number of iterations as that of the ChFSI method employing FP64 arithmetic. In contrast, the ChFSI method with FP32 and TF32 arithmetic fails to reach the same residual tolerance as that of the ChFSI method with FP64 arithmetic.
Input:
Chebyshev polynomial order $p$, estimates of the bounds of the eigenspectrum $\lambda_{max},\lambda_{min}$, estimate of the upper bound of the wanted spectrum $\lambda_{T}$ and the initial guess of eigenvectors $\textbf{X}^{(i)}$ and eigenvalues $\bm{\Lambda}^{(i)}$
Result: The filtered subspace $\textbf{Y}_{p}^{(i)}$
Temporary Variable: X, Y, $\textbf{R}_{\textbf{X}}$, $\textbf{R}_{\textbf{Y}}$,$\bm{\Lambda}_{\textbf{X}}$ and $\bm{\Lambda}_{\textbf{Y}}$
1 $e\leftarrow\frac{\lambda_{max}-\lambda_{T}}{2}$; $c\leftarrow\frac{\lambda_{max}+\lambda_{T}}{2}$;
2 $\sigma\leftarrow\frac{e}{\lambda_{min}-c}$;
3 $\sigma_{1}\leftarrow\sigma$;
4 $\gamma\leftarrow\frac{2}{\sigma_{1}}$;
5 $\textbf{X}\leftarrow\textbf{X}^{(i)}$;
6 $\textbf{Y}\leftarrow\textbf{A}\textbf{X}^{(i)}-\textbf{X}^{(i)}\bm{\Lambda}^{(i)}$;
7 $\textbf{R}_{\textbf{X}}\leftarrow 0$;
8 $\textbf{R}_{\textbf{Y}}\leftarrow\frac{2\sigma_{1}}{e}\textbf{Y}$;
9 $\bm{\Lambda}_{\textbf{X}}\leftarrow\textbf{I}$;
10 $\bm{\Lambda}_{\textbf{Y}}\leftarrow\frac{2\sigma_{1}}{e}\left(\bm{\Lambda}^{(i)}-c\textbf{I}\right)$;
11 for $k\leftarrow 2$ to $p$ do
12       $\sigma_{2}\leftarrow\frac{1}{\gamma-\sigma}$;
13       $\textbf{R}_{\textbf{X}}\leftarrow\frac{2\sigma_{2}}{e}\textbf{A}\textbf{R}_{\textbf{Y}}-\frac{2\sigma_{2}}{e}c\textbf{R}_{\textbf{Y}}-\sigma\sigma_{2}\textbf{R}_{\textbf{X}}+\frac{2\sigma_{2}}{e}\textbf{Y}\bm{\Lambda}_{\textbf{Y}}$;
14       $\bm{\Lambda}_{\textbf{X}}\leftarrow\frac{2\sigma_{2}}{e}\bm{\Lambda}_{\textbf{Y}}\bm{\Lambda}^{(i)}-\frac{2\sigma_{2}}{e}c\bm{\Lambda}_{\textbf{Y}}-\sigma\sigma_{2}\bm{\Lambda}_{\textbf{X}}$;
@@ 12       $\sigma_{2}\leftarrow\frac{1}{\gamma-\sigma}$;
16       swap$\left(\bm{\Lambda}_{\textbf{X}},\bm{\Lambda}_{\textbf{Y}}\right)$;
17       $\sigma=\sigma_{2}$;
18
[…]
@@ 17       $\sigma=\sigma_{2}$;
Algorithm 2 Reformulated ChFSI (R-ChFSI) Procedure for standard Hermitian eigenvalue problems
@@ 4.1.1 Symmetric Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations sampled at the origin (Gamma point) of the Brillouin zone for the material systems described in Table 1. This ensures that the resulting discretized equation is a real symmetric eigenvalue problem. The computation of the matrix multi-vector products with FP64 arithmetic is performed according to the methodology prescribed by [38, 55]. When using FP32 arithmetic to evaluate $\textbf{A}\textbf{R}_{\textbf{Y}}$ in Algorithm 2 of Algorithm 2, the data structures storing the matrix A and multi-vectors $\textbf{R}_{\textbf{X}}$ and $\textbf{R}_{\textbf{Y}}$ are in FP32 format and the BLAS calls used to perform $\textbf{A}\textbf{R}_{\textbf{Y}}$ are replaced with corresponding FP32 variants while the addition and scaling of multi-vectors is done using FP64 arithmetic. For TF32 arithmetic on NVIDIA GPUs we replace the cublasSgemm and its strided/batched variants with cublasGemmEx and its strided/batched variants with computeType set to CUBLAS_COMPUTE_32F_FAST_TF32. From Fig. 1, we observe that for all the benchmark systems summarized in Table 1, the accuracy achieved by the FP32 and TF32 variants for the R-ChFSI method is comparable to that achieved by the FP64 variant of the ChFSI method, whereas the FP32 and TF32 variants of the ChFSI method show significant reduction in achievable residual tolerance. We also note that the rate of convergence that is achieved by the FP32 and TF32 variants for the R-ChFSI method is comparable to that achieved by the FP64 variant of the ChFSI method. We also report the speedups achieved by the FP32 and TF32 variants of ChFSI and R-ChFSI over the FP64 variant of ChFSI in Fig. 2 on NVIDIA H100 GPU accelerators. We note that we obtain similar speedups for both the ChFSI and R-ChFSI methods allowing us to conclude that the additional overheads in the R-ChFSI method are negligible. We have achieved speedups of up to 1.75x for the TF32 variant of the R-ChFSI method over the baseline implementation of the ChFSI method with FP64 arithmetic. We also report the speedups achieved by the FP32 variant of ChFSI and R-ChFSI over the FP64 variant of ChFSI in Fig. 3 on Intel Xeon Platinum 8480C CPUs. We note that the speedups of the R-ChFSI method are slightly lower than ChFSI method due to the additional overheads in the R-ChFSI method. We achieve speedups of up to 2.1x for the FP32 variant of the R-ChFSI method over the baseline implementation of the ChFSI method with FP64 arithmetic.

### v3

@@ 3.3.2 Proposed residual-based Chebyshev filtering approach for generalized eigenproblem
$\implies$ $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq\zeta(\lVert\textbf{B}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,+\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}\rVert\,)$
We further note that using Lemma 3.11 we can write
$\lVert\textbf{D}^{-1}\textbf{B}\underline{\textbf{R}}_{k}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{k}^{(i)}+\textbf{D}^{-1}\textbf{B}\textbf{R}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+f_{k}\lVert\textbf{D}^{-1}\textbf{B}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,$
Thus we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq\zeta\lVert\textbf{A}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\zeta(f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\lVert\textbf{B}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,)$. Now, using Lemma 3.5 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=0$ and $\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\!\!=\!\lVert\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{B}\textbf{R}^{(i)}\rVert\,\!\!\leq\lvert\frac{\sigma_{1}}{e}\rvert\,\zeta\lVert\textbf{B}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,$, $h_{0}=\zeta\lVert\textbf{A}\rVert\,$ and $h_{1}^{(i)}=\zeta(f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\lVert\textbf{B}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,)\lVert\textbf{R}^{(i)}\rVert\,$ we obtain $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\zeta\tilde{\eta}_{k}\lVert\textbf{R}^{(i)}\rVert\,$ with $\tilde{\eta}_{k}=\lvert\frac{\sigma_{1}}{e}\rvert\,\lVert\textbf{B}\rVert\,g_{k-1}^{0}+(f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\lVert\textbf{B}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,)g_{k-1}^{1}$.
@@
(26) $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\zeta\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$
and if this inequality is satisfied for at $i=i_{0}$ then it holds for all $i>i_{0}$
Proof 3.26.
As a consequence of Lemmas 3.11, 3.23, and 3.15 we can now write $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,\zeta\tilde{\eta}_{p}\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ and in order to have $\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})<\angle(\mathcal{S}^{(i)},\mathcal{S})$ from Eq. 3 we require
(27) $\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\zeta\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$
@@ 4 Results and Discussion
In the following subsections, we focus on specific problem classes, starting with the results for standard eigenvalue problems, including the cases of real symmetric and complex Hermitian matrices, followed by generalized eigenvalue problems, where we assess the convergence behavior and the performance of the ChFSI and the R-ChFSI methods. To this end, we consider a set of three benchmark systems whose dimensions are summarized in Table 1. These eigenvalue problems are obtained from the finite-element discretization of the Kohn-Sham DFT equations for the periodic supercells of various sizes for body-centered cubic Molybdenum systems with a single vacancy. For all the benchmarks we utilize a Chebyshev polynomial degree of 75.
Table 1: Dimensions of the benchmark problems considered. These correspond to 3 material systems comprising $6\times 6\times 6$, $8\times 8\times 8$ and $10\times 10\times 10$ body-centered-cubic periodic supercells of Molybdenum with a single vacancy.
@@ 4.1 Standard Eigenvalue Problems
We now evaluate the accuracy and performance of the proposed R-ChFSI method as described in Algorithm 2 for solving the standard eigenvalue problems. We choose the eigenproblem encountered in using the higher-order finite-element discretization of Kohn-Sham DFT as outlined in [7, 28]. We begin by comparing the R-ChFSI method to the ChFSI method using FP32 and TF32 arithmetic. We demonstrate below that the R-ChFSI method with FP32 and TF32 arithmetic converges to the desired residual tolerance in a similar number of iterations as that of the ChFSI method employing FP64 arithmetic. In contrast, the ChFSI method with FP32 and TF32 arithmetic fails to reach the same residual tolerance as that of the ChFSI method with FP64 arithmetic.
Algorithm 2 R-ChFSI procedure for standard Hermitian eigenvalue problems
INPUTS: Chebyshev polynomial order $p$, estimates of the bounds of the eigenspectrum $\lambda_{max},\lambda_{min}$, estimate of the upper bound of the wanted spectrum $\lambda_{T}$ and the initial guess of eigenvectors $\textbf{X}^{(i)}$ and eigenvalues $\boldsymbol{\Lambda}^{(i)}$
OUTPUT: The filtered subspace $\textbf{Y}_{p}^{(i)}$
TEMPORARY VARIABLES:X, Y, $\textbf{R}_{\textbf{X}}$, $\textbf{R}_{\textbf{Y}}$,$\boldsymbol{\Lambda}_{\textbf{X}}$ and $\boldsymbol{\Lambda}_{\textbf{Y}}$
  $e\leftarrow\frac{\lambda_{max}-\lambda_{T}}{2}$; $c\leftarrow\frac{\lambda_{max}+\lambda_{T}}{2}$; $\sigma\leftarrow\frac{e}{\lambda_{min}-c}$; $\sigma_{1}\leftarrow\sigma$; $\gamma\leftarrow\frac{2}{\sigma_{1}}$
  $\textbf{X}\leftarrow\textbf{X}^{(i)}$; $\textbf{Y}\leftarrow\textbf{A}\textbf{X}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$
  $\textbf{R}_{\textbf{X}}\leftarrow 0$; $\textbf{R}_{\textbf{Y}}\leftarrow\frac{2\sigma_{1}}{e}\textbf{Y}$
  $\boldsymbol{\Lambda}_{\textbf{X}}\leftarrow\textbf{I}$; $\boldsymbol{\Lambda}_{\textbf{Y}}\leftarrow\frac{2\sigma_{1}}{e}\left(\boldsymbol{\Lambda}^{(i)}-c\textbf{I}\right)$
  for $k\leftarrow 2$ to $p$ do
     $\sigma_{2}\leftarrow\frac{1}{\gamma-\sigma}$
     $\textbf{R}_{\textbf{X}}\leftarrow\frac{2\sigma_{2}}{e}\textbf{A}\textbf{R}_{\textbf{Y}}-\frac{2\sigma_{2}}{e}c\textbf{R}_{\textbf{Y}}-\sigma\sigma_{2}\textbf{R}_{\textbf{X}}+\frac{2\sigma_{2}}{e}\textbf{Y}\boldsymbol{\Lambda}_{\textbf{Y}}$
     $\boldsymbol{\Lambda}_{\textbf{X}}\leftarrow\frac{2\sigma_{2}}{e}\boldsymbol{\Lambda}_{\textbf{Y}}\boldsymbol{\Lambda}^{(i)}-\frac{2\sigma_{2}}{e}c\boldsymbol{\Lambda}_{\textbf{Y}}-\sigma\sigma_{2}\boldsymbol{\Lambda}_{\textbf{X}}$
@@
  end for
  $\textbf{X}\leftarrow\textbf{R}_{\textbf{Y}}+\textbf{X}\boldsymbol{\Lambda}_{\textbf{Y}}$
@@ 4.1.1 Symmetric Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations sampled at the origin (Gamma point) of the Brillouin zone for the material systems described in Table 1. This ensures that the resulting discretized equation is a real symmetric eigenvalue problem. The computation of the matrix multi-vector products with FP64 arithmetic is performed according to the methodology prescribed by [7, 28]. When using FP32 arithmetic to evaluate $\textbf{A}\textbf{R}_{\textbf{Y}}$ in Algorithm 2, the data structures storing the matrix A and multi-vectors $\textbf{R}_{\textbf{X}}$ and $\textbf{R}_{\textbf{Y}}$ are in FP32 format and the BLAS calls used to perform $\textbf{A}\textbf{R}_{\textbf{Y}}$ are replaced with corresponding FP32 variants while the addition and scaling of multi-vectors is done using FP64 arithmetic. For TF32 arithmetic on NVIDIA GPUs we replace the cublasSgemm and its strided/batched variants with cublasGemmEx and its strided/batched variants with computeType set to CUBLAS_COMPUTE_32F_FAST_TF32.
@@
From Fig. 1, we observe that for all the benchmark systems summarized in Table 1, the accuracy achieved by the FP32 and TF32 variants for the R-ChFSI method is comparable to that achieved by the FP64 variant of the ChFSI method, whereas the FP32 and TF32 variants of the ChFSI method show significant reduction in achievable residual tolerance. We also note that the rate of convergence that is achieved by the FP32 and TF32 variants for the R-ChFSI method is comparable to that achieved by the FP64 variant of the ChFSI method.
@@
We also report the speedups achieved by the FP32 and TF32 variants of ChFSI and R-ChFSI over the FP64 variant of ChFSI in Fig. 2 on NVIDIA H100 GPU accelerators. We note that we obtain similar speedups for both the ChFSI and R-ChFSI methods allowing us to conclude that the additional overheads in the R-ChFSI method are negligible. We have achieved speedups of up to 1.75x for the TF32 variant of the R-ChFSI method over the baseline implementation of the ChFSI method with FP64 arithmetic.
@@
We also report the speedups achieved by the FP32 variant of ChFSI and R-ChFSI over the FP64 variant of ChFSI in Fig. 3 on Intel Xeon Platinum 8480C CPUs. We note that the speedups of the R-ChFSI method are slightly lower than ChFSI method due to the additional overheads in the R-ChFSI method. We achieve speedups of up to 2.1x for the FP32 variant of the R-ChFSI method over the baseline implementation of the ChFSI method with FP64 arithmetic.